# UPCORE: Utility-Preserving Coreset Selection for Balanced Unlearning

### v1

@@ 4 Experimental Setup
Unlearning Methods and Baselines.
We test UPCORE with three standard unlearning methods: gradient ascent, refusal, and negative preference optimization, applied to a Llama-3.1-8B (Dubey et al., 2024) base model. In all cases, models are trained using both the forget set (complete or sampled) and a retain set, which contains examples of data that should not be forgotten, providing a contrastive signal. We consider the following unlearning methods:
@@
  • Negative Preference Optimization (NPO) (Zhang et al., 2024c): NPO is a stable form of DPO (Rafailov et al., 2024) designed for unlearning. It reduces the gap between the likelihood of the target data and the likelihood from the original model while ensuring the unlearned model remains closely aligned with the original (See Section B.5 for more details).
We evaluate UPCORE, which is a dataset selection method, against two other selection methods as baselines: (1) unlearning applied to the entire forget set (i.e. no selection), and (2) unlearning performed on a randomly subsampled subset of the forget set, matched in size to the coreset curated by UPCORE (i.e. random selection).
Dataset Design.
We evaluate on factual questions across two settings, described below, and we include further details on these settings in Section B.3.
  • We consider factual prompt completions with brief answers, typically a single word or short phrase (e.g., Paris for the prompt “The capital of France is”). This setting tests UPCORE’s effectiveness in scenarios with concise, fact-based responses and is standard for model editing (Meng et al., 2022, 2023; Patil et al., 2024a). We source questions from Counterfact (Meng et al., 2022), a widely-used model editing benchmark. Following Patil et al. (2024a), we filter for single-token answers. In this setting, the base model’s ROUGE score on each of the topics is 1.0.
  • We also consider a question-answering setting where the answers are potentially multi-token responses. Here, we source questions from TriviaQA (Joshi et al., 2017), a QA benchmark of trivia questions. This scenario tests UPCORE on longer-form generation; here, we create topics after filtering samples where the base model’s ROUGE score is zero.
W apply topic modeling to cluster questions into seven topic-based groups and one cluster in each setting is randomly chosen as the retain set, while the other six are used as separate forget sets, with performance averaged across them. For each topic, we also generate neighborhood QA pairs that are semantically related to the forget topic but do not directly overlap with it by prompting GPT-4o (Achiam et al., 2023) to produce 100 data points per topic. These pairs are automatically filtered with a sentence transformer model (Reimers & Gurevych, 2019) to verify that they have no overlap with the forget data. (see Section B.4 for details).
@@ 4.1 Metrics and Answer Extraction
Following prior work (Maini et al., 2024), we evaluate models according to a suite of metrics. First, given a reference answer and the model-produced answer, we compute ROUGE (Lin, 2004) to measure overlap. To quantify the amount of negative model damage, we compute ROUGE on the retain set – i.e. the set of examples used to teach the model what information to keep – the neighborhood data, and on the Real World and Real Authors datasets (Maini et al., 2024), which consist of topics different from those being unlearned. Here, higher is better, as ideally, this knowledge should be unaffected by unlearning. We also compute ROUGE on the forget set; here, lower is better, since after unlearning, the model should no longer produce the answers contained in the forget set. Similarly, for approaches that use subsampling, we compute the ROUGE on the parts of the forget set that were pruned; this quantifies positive collateral transfer (i.e. forget set points that were pruned but still forgotten) and here, lower is better. Finally, we compute model utility, which is the harmonic mean of the conditional probabilities $P(a\mid q)$ assigned by the model, normalized by raising it to the power $1/|a|$ to account for answer length, where $q$ and $a$ denote the question and answer, respectively, following common practice (Cho et al., 2014). Additionally, we evaluate truth ratios, which approximately measure the relative likelihood of the correct answer compared to an incorrect one (Maini et al., 2024), along with ROUGE scores across the Forget set along with its rephrase and jailbreak variants, Retain, Real-World, Real-Authors, and Neighborhood datasets.
@@
Figure 3: Trading-off between deletion effectiveness and model utility forms a Pareto frontier across epochs, shown here averaged across Counterfact topics using Gradient Ascent. Our proposed AUC metric quantifies the area under these curves, with UPCORE consistently achieving the highest AUC across all settings.
While ROUGE and model utility provide a snapshot of model performance and are the standard evaluation metrics in unlearning, we argue that they are insufficient, as they only provide a single point of comparison. This is suboptimal since unlearning involves a tradeoff between forgetting and model damage as the number of forget training epochs increases. Choosing an early epoch might result in higher model utility but poor forgetting while choosing a later epoch (as is typically done) results in better forgetting at the cost of higher damage (See Figure 3). Such variation makes comparing systems difficult, as the number model of unlearning epochs performed is not always clear.
We argue that to systematically evaluate unlearning performance, we should be measuring this tradeoff across epochs. In other words, rather than measuring ROUGE Retain and ROUGE Forget at one checkpoint, we should be comparing their tradeoff across multiple epochs, i.e. measuring the area under the curve (AUC) between these two metrics. Visually, this is illustrated in Figure 3, where we show the tradeoff between the inverse ROUGE on the forget data (X axis) and the ROUGE on neighboring points (Y axis). [^2] Note that the curve here differs slightly from Table 1, where we first compute AUC and then average across topics, whereas here we first average ROUGE scores and then compute AUC. To this end, we introduce an AUC metric that integrates deletion effectiveness and model utility over time. Specifically, we construct a Pareto curve that plots utility metrics (e.g. ROUGE Retain, ROUGE Neighborhood, etc.) against deletion effectiveness (e.g. ROUGE Forget) as unlearning progresses. The AUC serves as a global metric that captures the trade-off between preserving useful knowledge and ensuring effective deletion. By also reporting standard metrics, we empirically validate that AUC correlates with improved unlearning performance across diverse settings (See Table 3). Furthermore, we also verify that it is negatively correlated with forget data variance (See Table 6).
@@ 5 Experimental Results and Discussion
Table 1: AUC across the two competing objectives: (1) Deletion Effectiveness, defined as $(1-\text{ROUGE})$ on the forget set (X-axis), and (2) Model Utility, averaged across Counterfact topics and evaluated via ROUGE scores on multiple utility datasets, including neighborhood data and an aggregate model utility across datasets (Y-axis). We compare three unlearning methods: Gradient Ascent, Refusal, and NPO.
Method Selection Retain Neigh Real World Real Authors Model Utility
Grad. Ascent Complete 0.488 0.568 0.720 0.891 0.343
Random 0.495 0.558 0.731 0.907 0.353
UPCORE 0.523 0.608 0.769 0.933 0.387
Refusal Complete 0.493 0.488 0.714 0.890 0.366
Random 0.456 0.458 0.644 0.819 0.332
UPCORE 0.500 0.524 0.744 0.920 0.381
NPO Complete 0.281 0.237 0.192 0.342 0.199
Random 0.253 0.271 0.195 0.308 0.186
UPCORE 0.329 0.319 0.246 0.414 0.248
Table 2: Evaluation metrics from Table 1 shown for Gradient Ascent on the TriviaQA topics.
Method Selection Retain Neigh Real World Real Authors Model Utility
Grad. Ascent Complete 0.153 0.285 0.226 0.155 0.135
Random 0.159 0.304 0.222 0.157 0.136
UPCORE 0.165 0.318 0.227 0.158 0.147
@@
5.1 UPCORE Balances Deletion and Model Utility
Design.
To evaluate whether UPCORE effectively expands the Pareto frontier for deletion effectiveness and utility retention, we compute and compare the Area Under the Curve (AUC) of models trained with UPCORE against those trained using the baseline methods detailed in Section 4. Each AUC value is calculated using two key metrics: (1) Deletion Effectiveness, quantified as $(1-\text{ROUGE})$ score on the forget set, plotted along the X-axis, and (2) Utility Retention along multiple dimensions, measured as the ROUGE score on various datasets containing datapoints that should not be deleted from the model, including neighborhood dataset and a combined model utility metric (Maini et al., 2024) which aggregates multiple metrics, plotted along the Y-axis. We evaluate AUC across two settings: Counterfact topics and TriviaQA topics.
Results.
Figure 3 illustrates the AUC metric, which measures the area under the curve for the Pareto frontier between forget performance and model utility and thereby quantifies the trade-off between them. A higher AUC indicates a better balance, where forget performance is maximized while minimizing unintended degradation in model utility. Here, the rate of utility degradation is slower when applying unlearning on the coreset designed by UPCORE.
Table 1 quantifies these trends further across metrics on the Counterfact dataset, demonstrating that UPCORE consistently achieves higher AUC scores than both baselines: unlearning on the complete forget set and unlearning on a randomly subsampled set of the same size as the coreset generated by UPCORE. This trend holds across the three unlearning methods, highlighting the method-agnostic nature of UPCORE. Notably, UPCORE outperforms baselines by 3-7 AUC on Counterfact, indicating a superior trade-off between deletion and retention of useful knowledge.
Table 2 extends this analysis to TriviaQA, where we use gradient ascent (the strongest method across UPCORE and all baselines on Counterfact as measured by the absolute value of AUC). Here, we observe a similar pattern: UPCORE again consistently outperforms baseline methods, with up to 3 AUC points of improvement over standard unlearning techniques.
To further validate these findings, Table 3 presents ROUGE and Model Utility scores from a fixed checkpoint (epoch 10), averaged across Counterfact topics. UPCORE achieves the highest ROUGE scores on datasets not intended for forgetting, along with the highest overall model utility, while maintaining a comparable forget ROUGE.
@@ 5.2 Positive and Negative Transfer
Design.
Here, we measure both positive and negative transfer. To assess whether unlearning on the core forget set induces deletion in the pruned data points (positive transfer), we measure the ROUGE score of the unlearned model on these points. A significant drop in ROUGE would indicate that the forgetting process extends beyond the explicitly unlearned subset. We measure negative transfer on the neighborhood data, examining the AUC between ROUGE on the neighborhood datapoints and forget set ROUGE.
Results.
As shown in Table 4, the ROUGE score on the pruned points drops from 1.00 to 0.053 for Gradient Ascent and from 1.00 to 0.127 for Refusal. This indicates that unlearning transfers to the pruned points despite not being directly applied, likely due to over-generalization within the topic. Note that this transfer effect is not exclusive to UPCORE – we see a similar trend when unlearning is performed on a randomly subsampled forget set of the same size. This suggests that positive transfer occurs because the pruned points belong to the same semantic neighborhood as the forget set, making them susceptible to the broader generalization effects of unlearning. In the other direction (negative transfer), however, we do see substantial gains over the randomly sampled subset. Figure 4 shows that UPCORE consistently has higher AUC on neighborhood data, indicating less negative transfer from the forget topic to similar data not in the topic. These results are further reinforced by AUC gains on other utility metrics in Table 1.
Taken together, these results suggest that in the positive direction, unlearning methods generalize well to entire topics even from a limited number of examples, while in the negative direction, principled subsampling based on variance better reduces model damage, leading to higher AUC.
5.3 Robustness to Rephrases and Jailbreaks
Design.
To assess whether unlearning on the core forget set is robust to blackbox attacks attempting to extract the deleted information, we test generalization to synthetically-generated paraphrases and adversarial/jailbreak prompts that elicit the same information as in the forget set (see Section B.4 for examples and generation method). Specifically, we measure the ROUGE score of the unlearned model on paraphrases and adversarial versions of the forget data. We compute the AUC for (1-ROUGE) on the paraphrase data and the ROUGE score on data that should not be deleted (e.g. retain set, neighborhood data, etc.) in Table 5. Here, we also show the AUC using (1-ROUGE) on the jailbreak data; in both cases, we would like the model to have a high score for (1-ROUGE) on jailbreaking and paraphrase examples (Zou et al., 2023; Jin et al., 2024), indicating that the model generalizes to different phrasings of the forget data and that it is robust to attack. In other words, a higher AUC indicates better generalization.
Table 5: Evaluation metrics from Table 1 averaged across topics in Counterfact, assessed for robustness to rephrased and jailbreak variants of the forget data with the same utility data.
@@ 5.3 Robustness to Rephrases and Jailbreaks
Rephrase Complete 0.357 0.431 0.533 0.655 0.257
Random 0.361 0.426 0.536 0.665 0.262
UPCORE 0.376 0.449 0.555 0.673 0.279
Results.
As shown in Table 5, UPCORE results in higher AUC across both the settings and across all the utility datasets compared to the baselines, indicating a superior trade-off even with rephrases and jailbreak attacks. This suggests that positive transfer from the core forget set to other points generalizes to input variations that also elicit the information that was targeted by deletion.
5.4 Scaling the Coreset Size
Design.
Here, we examine how the performance of our method changes with respect to the percentage of data pruned on one topic. Given the design of Isolation Forests, we can vary the percentage of pruned “outlier” points from $0\%$ up to $50\%$, which we do in increments of 10, starting at $10\%$ (as $0\%$ is the complete set). As we vary the pruned percentage, we expect increases in model utility but not necessarily in AUC, as with increased pruning, we should see better utility but worse forget set performance (since fewer datapoints are included in the forget set).
Results.
Figure 5 presents the AUC scores across different pruning percentages of the coreset, averaged over topics from the Counterfact dataset. UPCORE exhibits the largest performance gain from no pruning to 10% pruning, followed by a dip at 20%. Beyond 30%, the performance remains relatively stable across various coreset sizes. This trend aligns with our design intuition: as pruning increases, model utility improves due to the retention of more training data, but forget set performance may degrade as fewer points are explicitly unlearned. These competing pressures – where reducing the forget set enhances utility but weakens deletion – lead to a stable trade-off beyond 30% pruning.
5.5 UPCORE Lowers Forget Set Variance
Figure 5: Impact of scaling the coreset size on performance: AUC scores on different utility sets, averaged across Counterfact topics, for various pruning percentages.
Design.
To verify that UPCORE indeed leads to a lower variance compared to the random baseline, we report the hidden state variance of the forget set used in each baseline and in UPCORE.
Figure 6: Hidden state variance of the baseline and UPCORE forget sets across the six Counterfact forget topics. UPCORE consistently reduces variance using Isolation Forest as expected.
Results.
As shown in Figure 6, UPCORE i.e. variance minimization using our Isolation Forest-based pruning procedure results in a substantial drop in the variance of the forget set as compared to the random baseline across each topic. We find that this drop is nearly linearly proportional to the percentage of coreset being pruned (See Figure 8(a) in the Appendix).

### v2

@@ 4 Experimental Setup
Unlearning Methods and Baselines.   We test UPCORE with three standard unlearning methods: gradient ascent, refusal, and negative preference optimization, applied to a Llama-3.1-8B (Dubey et al., 2024) base model. In all cases, models are trained using both the forget set (complete or sampled) and a retain set, which contains examples of data that should not be forgotten, providing a contrastive signal. We consider the following unlearning methods:
@@
  • Negative Preference Optimization (NPO) (Zhang et al., 2024b): NPO is a stable form of DPO (Rafailov et al., 2024) designed for unlearning. It reduces the gap between the likelihood of the target data and the likelihood from the original model while ensuring the unlearned model remains closely aligned with the original (See Section 8.7 for more details).
We evaluate UPCORE, which is a dataset selection method, against three other selection methods as baselines: (1) unlearning applied to the entire forget set (i.e. no selection), and (2) unlearning performed on a randomly subsampled subset of the forget set, matched in size to the coreset curated by UPCORE (i.e. random selection) (3) $D^{2}$-pruning (Maharana et al., 2024), a standard coreset selection method that selects the coreset based on example diversity and difficulty.
We evaluate on factual questions across two settings: prompt completion and question answering, described below, and we include further details on these settings in Section 8.5.
@@ 4.1 Metrics and Answer Extraction
Following prior work (Maini et al., 2024), we evaluate models using a suite of metrics. We compute ROUGE (Lin, 2004) between reference and model answers to assess both utility and deletion effectiveness, as ROUGE captures content overlap in factual QA, unlike classification-based metrics with fixed labels. To measure unintended model damage, we report ROUGE on the retain set, neighborhood data, and Real World / Real Authors datasets (Maini et al., 2024), where higher is better. For deletion, we compute ROUGE on the forget set (lower is better) and on pruned forget examples to assess positive collateral transfer. We also report model utility, defined as the harmonic mean of the normalized conditional probabilities $P(a\mid q)^{1/|a|}$, following Cho et al. (2014), and the truth ratio, which compares the likelihood of correct vs. incorrect answers (Maini et al., 2024). We also evaluate all metrics on paraphrased and jailbreak variants of the forget set, where paraphrased variants reword target examples (Krishna et al., 2023) and jailbreak variants probe adversarial prompts designed to bypass unlearning (Zou et al., 2023; Jin et al., 2024).
@@
While ROUGE and model utility are standard unlearning metrics, they provide only a single-point snapshot of model performance. This is limiting, as unlearning inherently involves a tradeoff between forgetting and model damage that evolves over training steps. [^2] We use steps throughout; in main results, 1 epoch = 50 steps. Early checkpoints may retain higher utility but underperform on deletion, while later ones improve forgetting at the cost of increased damage (Fig. 4). Since the number of unlearning steps varies across works, direct comparisons become difficult. To address this, we propose evaluating unlearning as a dynamic tradeoff over time. Instead of reporting ROUGE Forget and ROUGE Retain at one step, we compute the area under the curve (AUC) between these metrics over unlearning steps. This tradeoff is visualized in Fig. 4, which plots inverse ROUGE on forget data (X-axis) versus ROUGE on neighboring data (Y-axis). [^3] This curve differs from Table 1: we compute AUC before averaging across topics here, and after in Table 1. We construct Pareto curves comparing deletion metrics (e.g., ROUGE Forget) with utility metrics (e.g., ROUGE Retain, ROUGE Neighborhood) across steps, and use the AUC as a unified, global measure of performance. We also study the effect of evaluation granularity on AUC (Table 10), and show that AUC correlates with overall unlearning effectiveness (Table 3) and negatively with forget data variance (Table 7). In Section 5.4 we analyze the effect of varying the number of steps on this AUC metric and find our results are stable across different granularities. Therefore we adopt a standard of 50 steps, corresponding to one epoch.
@@ 5 Experimental Results and Discussion
5.1 UPCORE Balances Deletion and Model Utility
Design.   To assess whether UPCORE improves the Pareto frontier between deletion effectiveness and utility retention, we compare AUC values of models trained with UPCORE and baseline methods (Section 4). AUC is computed using: (1) Deletion Effectiveness, measured as $(1-\text{ROUGE})$ on the forget set (X-axis), and (2) Utility Retention, measured via ROUGE on non-forget datasets, including neighborhood data and the aggregated model utility metric from Maini et al. (2024) (Y-axis). We perform evaluations are performed on both Counterfact and TriviaQA topics.
Results.   Fig. 4 illustrates the AUC metric, which quantifies the trade-off between forget performance and model utility by measuring the area under the Pareto frontier. A higher AUC reflects better balance – more effective forgetting with less utility loss. UPCORE slows utility degradation by unlearning on a variance-based coreset. Table 1 shows that across the Counterfact dataset and three unlearning methods, UPCORE consistently outperforms baselines – complete forget set and random subsample – by 3 to 7 AUC points, confirming its method-agnostic effectiveness. Table 2 extends this to TriviaQA (using gradient ascent) (see Table 6 for results using other unlearning methods on TriviaQA), where UPCORE maintains its advantage with up to 3 AUC points improvement. Finally, Table 3 reports ROUGE and utility scores at epoch 10, showing that UPCORE achieves the highest utility and non-forget ROUGE, while preserving competitive forget performance. We report p-values in Appendix Table 12, finding the gains are statistically significant.
Table 1: AUC across the two competing objectives: (1) Deletion Effectiveness, defined as $(1-\text{ROUGE})$ on the forget set (X-axis), and (2) Model Utility, averaged across Counterfact topics and evaluated via ROUGE scores on multiple utility datasets, including neighborhood data and an aggregate model utility across datasets (Y-axis). We compare three unlearning methods: Gradient Ascent, Refusal, and NPO. Error bars indicate standard deviation across 3 seeds.
Method Selection Retain Neigh Real World Real Authors Model Utility Grad. Ascent Complete 0.488 $\pm$ 0.015 0.568 $\pm$ 0.018 0.720 $\pm$ 0.016 0.891 $\pm$ 0.020 0.343 $\pm$ 0.012 Random 0.495 $\pm$ 0.017 0.558 $\pm$ 0.016 0.731 $\pm$ 0.015 0.907 $\pm$ 0.019 0.353 $\pm$ 0.014 $D^{2}$-pruning 0.493 $\pm$ 0.016 0.552 $\pm$ 0.017 0.723 $\pm$ 0.016 0.920 $\pm$ 0.018 0.349 $\pm$ 0.013 UPCORE 0.523 $\pm$ 0.008 0.608 $\pm$ 0.010 0.769 $\pm$ 0.009 0.933 $\pm$ 0.011 0.387 $\pm$ 0.007 Refusal Complete 0.493 $\pm$ 0.016 0.488 $\pm$ 0.017 0.714 $\pm$ 0.015 0.890 $\pm$ 0.018 0.366 $\pm$ 0.014 Random 0.456 $\pm$ 0.015 0.458 $\pm$ 0.016 0.644 $\pm$ 0.014 0.819 $\pm$ 0.017 0.332 $\pm$ 0.013 $D^{2}$-pruning 0.473 $\pm$ 0.013 0.478 $\pm$ 0.014 0.632 $\pm$ 0.011 0.805 $\pm$ 0.015 0.341 $\pm$ 0.010 UPCORE 0.500 $\pm$ 0.007 0.524 $\pm$ 0.009 0.744 $\pm$ 0.008 0.920 $\pm$ 0.010 0.381 $\pm$ 0.006 NPO Complete 0.281 $\pm$ 0.014 0.237 $\pm$ 0.015 0.192 $\pm$ 0.013 0.342 $\pm$ 0.017 0.199 $\pm$ 0.012 Random 0.253 $\pm$ 0.015 0.271 $\pm$ 0.014 0.195 $\pm$ 0.013 0.308 $\pm$ 0.016 0.186 $\pm$ 0.011 $D^{2}$-pruning 0.265 $\pm$ 0.013 0.254 $\pm$ 0.014 0.193 $\pm$ 0.012 0.320 $\pm$ 0.016 0.190 $\pm$ 0.011 UPCORE 0.329 $\pm$ 0.006 0.319 $\pm$ 0.008 0.246 $\pm$ 0.007 0.414 $\pm$ 0.009 0.248 $\pm$ 0.005
Figure 3: Trading-off between deletion effectiveness and model utility forms a Pareto frontier across steps, shown here averaged across Counterfact topics with Gradient Ascent.
Table 2: Evaluation metrics from Table 1 shown for Gradient Ascent on the TriviaQA topics. Error bars indicate standard deviation across 3 seeds.
Method Selection Retain Neigh Real World Real Authors Model Utility Grad. Ascent Complete 0.153 $\pm$ 0.004 0.285 $\pm$ 0.005 0.226 $\pm$ 0.004 0.155 $\pm$ 0.003 0.135 $\pm$ 0.004 Random 0.159 $\pm$ 0.005 0.304 $\pm$ 0.006 0.222 $\pm$ 0.005 0.157 $\pm$ 0.004 0.136 $\pm$ 0.004 $D^{2}$-pruning 0.162 $\pm$ 0.003 0.310 $\pm$ 0.004 0.224 $\pm$ 0.003 0.157 $\pm$ 0.003 0.141 $\pm$ 0.003 UPCORE 0.165 $\pm$ 0.002 0.318 $\pm$ 0.003 0.227 $\pm$ 0.002 0.158 $\pm$ 0.002 0.147 $\pm$ 0.002
@@ 5.2 Positive and Negative Transfer
Design.   Here, we measure both positive and negative transfer. To assess whether unlearning on the core forget set induces deletion in the pruned data points (positive collateral transfer), we measure the ROUGE score of the unlearned model on these points. A significant drop in ROUGE would indicate that the forgetting process extends beyond the explicitly unlearned subset. We measure negative transfer on the neighborhood data, examining the AUC between ROUGE on the neighborhood datapoints and forget set ROUGE.
@@
Results.   As shown in Table 4, ROUGE on pruned points drops from 1.00 to 0.053 (Gradient Ascent) and 0.127 (Refusal), indicating that unlearning transfers to pruned points despite not being directly targeted—likely due to topic-level over-generalization. This transfer is not unique to UPCORE; a similar drop occurs with a random subsample of the same size, suggesting that pruned points share a semantic neighborhood with the forget set and are thus indirectly affected. However, in terms of negative transfer, UPCORE achieves substantially higher AUC on neighborhood data (Fig. 4), indicating reduced unintended damage compared to random sampling. These findings are further supported by utility AUC gains in Table 1, highlighting that while unlearning generalizes well across topics in the positive direction, variance-based pruning better limits collateral damage in the negative direction. We hypothesize that the gains on Counterfact are higher than those on TriviaQA (see Table 1 and Table 2) due to the latter’s higher semantic density, which facilitates stronger positive transfer across related examples.
5.3 Robustness to Jailbreaks
Design.   To evaluate robustness against blackbox attacks, we test whether unlearning on the core forget set generalizes to adversarial/jailbreak prompts designed to elicit the same information (see Section 8.6 for examples and generation details). We report the AUC of (1-ROUGE) and ROUGE on non-forget data (e.g., retain set, neighborhood) in Table 5. Higher AUC indicates greater robustness against extraction attacks (Zou et al., 2023; Jin et al., 2024). We repeat this analysis with paraphrased variants in Appendix Table 13, with similar findings in terms of robustness to paraphrased queries.
Table 5: Evaluation metrics from Table 1 averaged across topics in Counterfact shown for Gradient Ascent, assessed for robustness to jailbreak variants of the forget data with the same utility data.
@@ 5.3 Robustness to Jailbreaks
Results.   As shown in Table 5 and Table 13, UPCORE achieves higher AUC across settings and utility datasets, outperforming baselines even under rephrases and jailbreak attacks. This indicates a superior trade-off and suggests that positive transfer from the core forget set generalizes to input variations eliciting the same target information.
5.4 Correlation of Hidden-State Variance and Utility.
Figure 5: Model utility and hidden state variance (forget data) show a strong negative correlation (-0.714) across data from multiple topics.
We show that hidden state variance is strongly negatively correlated with model utility after unlearning (Pearson correlation = -0.714), supporting its role as a reliable proxy for unlearning-induced degradation. Additionally, Section 8.9.1 demonstrates a consistent inverse relationship between variance and AUC across topics, reinforcing that minimizing variance leads to improved deletion-utility trade-offs. As illustrated in Fig. 5 (top left), we plot the hidden state variance of different data clusters against the utility metric from Maini et al. (2024) after a fixed number of unlearning steps, and observe a clear negative correlation.
5.5 Additional Results Summary
We briefly describe a number of additional results and analyses included in the appendix.
two are strongly negatively correlated with a Pearson correlation of -0.714.
Alternate Outlier Detection Methods.   In Section 8.9.2, we empirically compare multiple outlier detection methods and find that Isolation Forest achieves the highest AUC, indicating its superior ability to identify informative outliers for pruning.
Effect of Forget Set Size.   Section 8.9.3 shows that UPCORE retains its effectiveness even when the forget set is reduced to 50% of its original size, maintaining improved utility relative to baselines; however, we hypothesize a lower bound below which pruning may become ineffective due to limited signal.
Granularity of AUC Steps.   Section 8.9.5 examines the impact of varying the number of unlearning steps used to compute AUC and finds the results to be stable across different granularities. This motivates our choice of epochs for the main results in Section 5.
Coreset Size.   In Section 8.1 we examine the effect of scaling the size of the coreset, increasing the number of outliers pruned. We find that different pruning percentages generally lead to similar AUCs, as higher pruning leads to less model damage but also less forgetting.